# Scalable Multi-view Clustering via Explicit Kernel Features Maps

### v1

@@ Abstract
A growing awareness of multi-view learning as an important component in data science and machine learning is a consequence of the increasing prevalence of multiple views in real-world applications, especially in the context of networks. In this paper we introduce a new scalability framework for multi-view subspace clustering. An efficient optimization strategy is proposed, leveraging kernel feature maps to reduce the computational burden while maintaining good clustering performance. The scalability of the algorithm means that it can be applied to large-scale datasets, including those with millions of data points, using a standard machine, in a few minutes. We conduct extensive experiments on real-world benchmark networks of various sizes in order to evaluate the performance of our algorithm against state-of-the-art multi-view subspace clustering methods and attributed-network multi-view approaches.
@@ 1 Introduction
The explosive growth of data from diverse sources, such as social media, sensor networks, and online platforms, has led to the emergence of complex high-dimensional datasets. These datasets often contain multiple views of the same underlying data, each capturing different aspects or perspectives. Traditional clustering algorithms, designed for single-view data, face significant challenges in effectively capturing the intricate relationships and structures extending across multiple views. In this context, multi-view clustering, which integrates information from multiple views in order to enhance clustering performance, has emerged as a powerful paradigm. It has been successfully applied to various applications including computer vision [28, 25], natural language processing [12, 16], and bioinformatics [20, 23, 21]. For a more in-depth discussion, please refer to [1].
@@
In this work we are specifically interested in multi-view subspace clustering [7, 22, 2]. By exploring the latent subspaces within each view and leveraging the complementary nature of different views, multi-view subspace clustering algorithms can reveal hidden structures that may not be apparent when considering individual views in isolation. This not only leads to more accurate clustering results, but also enables a deeper understanding of the underlying data. Specifically, multi-view subspace clustering for attributed network data comprising interconnected entities with associated attributes offers a powerful framework for analyzing complex datasets, and for this reason it has gained significant attention in recent literature [4, 13, 14, 17]. One prominent application of multi-view clustering for attributed network data is in social network analysis, given that social networks often exhibit rich attributes relating to individuals or groups – demographic information, interests, affiliations, and so on [4, 9]. Another application is in biological research, where multi-omics data can be integrated using multiple graph convolutional networks [24]. Finally, recommender systems often rely on multiple sources of information, such as user preferences, item attributes, and social connections [11, 27, 19].
@@
The computational complexity of multi-view subspace clustering has hindered its widespread adoption in real-world applications, even though some efficient approaches have been proposed. These can involve, for example, anchor-based techniques for reducing the size of matrices used in optimization [10, 22]. However, tests that we performed show that existing methods often fail to achieve the desired level of efficiency and clustering performance in the case of large-scale multi-view datasets, and consequently that further advances in this area need to be made to unlock the full potential of multi-view subspace clustering.
In this paper, rather than using anchor-based methods, we tackle the scalable multi-view subspace clustering problem via a novel algorithm that is able to improve on the clustering accuracy and computational efficiency of existing approaches. Our algorithm leverages low-rank subspace clustering and properties of kernel feature maps in order to capture complex relationships between views and improve the robustness of clustering results. Similarly, to some state-of-the-art models, our approach scales linearly with the number of inputs, but it scales more favorably with other factors such as the number of dimensions and clusters. Additionally, it harnesses the inherent parallelism present in the multi-view setting such that large-scale datasets can be processed efficiently. In summary, the key contributions of this paper are threefold:
@@
    Second, we propose an efficient optimization strategy that scales well with very large-scale datasets. By leveraging the properties of kernel features maps, our algorithm significantly reduces the computational burden while maintaining state-of-the-art clustering performance. This scalability makes our approach suitable for handling real-world applications, since it is able to scale to datasets with millions of data points.
@@
    Third, we conduct extensive experiments, including statistical significance testing, on real-world benchmark networks of different scales and degrees of overlap, to evaluate our algorithm against state-of-the-art multi-view subspace clustering methods and attributed-network-specific multi-view approaches. Our experiments are entirely reproducible, ensuring transparency and enabling others to validate our results.
The remainder of this paper is structured as follows. In section 2 we discuss related works. Section 3 introduces the necessary background to develop our approach. In section 4 we develop our proposed method and explore the algorithm and its complexity. Section 5 contains a detailed description of our experiments. Finally, we give our conclusion in section 6. A schematic representation of our approach is shown in Figure 1.
Figure 1: Schematic diagram: the blue section represents the operations that are explicitly performed in our approach; the green section represents section represents what is transpiring implicitly (operations in the same column represent the same step). Given a set of inputs, we apply an SVD to obtain the left singular values (equivalent to implicitly computing a coefficient matrix using low-rank subspace clustering). We then apply a nonnegative feature map (equivalent to implicitly computing an affinity matrix). We then weigh and concatenate the obtained embeddings of each view (equivalent to summing the affinity matrices into a single consensus one). Finally, we perform an SVD followed by a k-Means to obtain a consensus partition (equivalent to performing spectral clustering on the consensus affinity matrix)
@@ 2 Related Work
This section discusses prior works relating to scalable multi-view subspace clustering. We also look at graph multi-view clustering, since we use attributed networks to evaluate performance.
@@ 2.1 Scalable Multi-view Subspace Clustering
In [14] Large-Scale Multi-View Subspace Clustering in Linear Time (LMVSC) [10] was introduced. This method, where smaller graphs are created for each view and then integrated in order to enable spectral clustering on a reduced graph size, derives from the concept of anchor graphs. The Scalable Multi-view Subspace Clustering with Unified Anchors (SMVSC) [22] approach integrates anchor learning and graph construction into a unified optimization framework, resulting in a more accurate representation of latent data distribution by learned anchors and a more discriminative clustering structure. In [2], an efficient orthogonal multi-view subspace clustering (OMSC) approach was proposed, featuring joint anchor learning, graph construction and partitioning. Recently, [29] developed a Flexible and Diverse Anchor Graph Fusion for Scalable Multi-view Clustering (FDAGF) that introduces a fusion strategy for multi-size anchor graphs.
@@ 2.2 Graph Multi-view Clustering
In [17], a multi-view contrastive graph clustering (MCGC) method was proposed, designed to learn a consensus graph with contrastive regularization over the learned graph matrix. [14] introduced a novel multi-view attributed graph clustering (MAGC) framework, which incorporates node attributes and graphs. It is an approach that stands out in three important respects: utilizing graph filtering for smooth node representation, rather than instead deep neural networks; learning a consensus graph from data to address noise and incompleteness in the original graph; and exploring high-order relations through a flexible regularizer design. The MvAGC model [13] is a simple yet effective approach for Multi-view Attributed Graph Clustering. It applies a graph filter to smooth the features without relying on neural network parameter learning. A reduced computational complexity is achieved through a novel anchor point selection strategy. In addition, a new regularizer is introduced to explore high-order neighborhood information. [6] introduced (LMGEC), a generic linear model for performing multi-view attributed graph representation learning and clustering simultaneously. It uses a simple weighing scheme based on inertia and a single-hop neighborhood propagation.
@@ 3 Preliminaries
In what follows, matrices are denoted in boldface uppercase letters and vectors in boldface lowercase letters. $[n]$ corresponds to the integer set $\{1,\ldots,n\}$. Tr is the trace operator. $\oplus$ denotes column-wise concatenation.
@@ 3.1 Subspace Clustering
Given a data matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ and a desired number of clusters $k$, most subspace clustering problems have the following form:
$\min_{\mathbf{C}}\quad$ $\|\mathbf{X}-\mathbf{C}\mathbf{X}\|^{2}\,+\,\Omega(\mathbf{C})\quad\text{s.t.}\quad\mathbf{C}\in\mathcal{C}$ (1)
where $\mathbf{C}$ is called a coefficient matrix, $\Omega$ is a regularization function and $\mathcal{C}$ is the set of feasible solutions. Once a solution $\mathbf{C}$ is found, a subspace affinity graph represented using its adjacency matrix $\mathbf{W}$ is constructed, typically as $\mathbf{W}=(|\mathbf{C}|+|\mathbf{C}^{\top}|)/2$. Finally, a partition of $k$ groups is obtained via spectral clustering on $\mathbf{W}$. Different values of $\Omega$ and $\mathcal{C}$ can call for different subspace clustering techniques: these might, for example, be low-rank approaches [15] or sparse approaches [3].
$\min_{\mathbf{C}\in\mathcal{C}}\quad$ $\|\mathbf{X}-\mathbf{C}\mathbf{X}\|^{2}\,+\,\|\mathbf{C}\|_{*}\quad\text{s.t.}\quad\mathbf{C}=\mathbf{C}^{\top}$ (2)
Note that all norms here refer to the Frobenius norm, unless otherwise specified.
@@ 3.2 Scalable Subspace Clustering
In [5], the authors proposed a scalable low-rank subspace clustering model. Specifically, the problem they considered is:
$\min_{\mathbf{U}}\quad$ $\|\mathbf{X}-\mathbf{U}\mathbf{U}^{\top}\mathbf{X}\|^{2}\quad\text{s.t.}\quad\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}.$ (3)
To address the problem they use nonnegative kernel maps to create an affinity matrix $\mathbf{W}=\Phi(\mathbf{U})\Phi(\mathbf{U})^{\top}.$ The advantage of computing the affinity matrix in this manner is that spectral partitioning can be obtained via singular values decomposition on $\Phi(\mathbf{U})$ instead of performing eigendecomposition on $\mathbf{W}$, leading to substantial gains in spatial and computational efficiencies.
@@ 3.3 Multi-view Subspace Clustering
$\min_{\mathbf{F},\mathbf{C}_{1},\ldots,\mathbf{C}_{V}}\quad$ $\sum_{v}\|\mathbf{X}_{v}-\mathbf{C}_{v}\mathbf{X}_{v}\|^{2}+\texttt{Tr}(\mathbf{F}^{\top}\mathbf{C}_{v}\mathbf{F})$ (4)
such that $\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1}$
$\mathbf{C}_{1},\ldots,\mathbf{C}_{V}\in\mathcal{C},$
where $\mathbf{C}_{v}$ is the coefficient matrix of the $v$-th view and $\mathbf{F}$ is a a consensus partition matrix. However, although the block structures in different view-specific coefficient matrices are similar, the magnitudes and the signs of the entries in each $C_{v}$ can vary greatly. To solve this issue, [7] proposed using the affinity matrices $\mathbf{W}_{v}$ rather than the coefficient matrices $\mathbf{C}_{v}$. We generalize their formulation by dropping the specific constraints that the authors used. This results in the following generic multi-view subspace clustering problem:
$\min_{\mathbf{F},\mathbf{C}_{1},..,\mathbf{C}_{V}}$ $\quad\sum_{v}\|\mathbf{X}_{v}-\mathbf{C}_{v}\mathbf{X}_{v}\|^{2}\quad+\texttt{Tr}\left(\mathbf{F}^{\top}\left(\mathbf{I}-\tfrac{|\mathbf{C}_{v}|+|\mathbf{C}_{v}^{\top}|}{2}\right)\mathbf{F}\right)$ (5)
such that $\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1},\mathbf{C}_{1},\ldots,\mathbf{C}_{V}\in\mathcal{C}.$
@@
4 Proposed Methodology
We consider the multi-view framework presented in problem 5 and like in 3 we introduce into it low-rank constraints, also using the same affinity matrix. However, here we are considering a symmetrically normalized affinity matrix rather than the unnormalized one. We obtain the following problem:
$\min_{\mathbf{F},\mathbf{U}_{1},\ldots,\mathbf{U}_{V}}\quad$ $\sum_{v}\|\mathbf{X}_{v}-\mathbf{U}_{v}\mathbf{U}_{v}^{\top}\mathbf{X}_{v}\|^{2}+\ \texttt{Tr}(\mathbf{F}^{\top}(\mathbf{I}-\mathbf{W}_{v})\mathbf{F})$ (6)
such that $\mathbf{W}_{v}=\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})\Phi(\mathbf{U}_{v})^{\top}\mathbf{D}_{v}^{-\frac{1}{2}}$
$\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1}$
where $\mathbf{U}\in\mathbb{R}^{n\times f}$ is a low-rank matrix, and $\mathbf{W}_{v}$ is the normalized affinity matrix of view $v$. If we were to solve this problem directly, the consensus partition $\mathbf{F}$ would be obtained via spectral clustering on

### v2

@@ Abstract
The proliferation of high-dimensional data from sources such as social media, sensor networks, and online platforms has created new challenges for clustering algorithms. Multi-view clustering, which integrates complementary information from multiple data perspectives, has emerged as a powerful solution. However, existing methods often struggle with scalability and efficiency, particularly on large attributed networks. In this work, we address these limitations by leveraging explicit kernel feature maps and a non-iterative optimization strategy, enabling efficient and accurate clustering on datasets with millions of points.
keywords:
multi-view clustering, community detection, graph clustering, subspace clustering
Figure 1: Schematic diagram: the black section represents the operations that are explicitly performed in our approach; the green section represents what is transpiring implicitly (operations in the same column represent the same step). Given a set of inputs, we apply an singular value decomposition (SVD) to obtain the left singular values (equivalent to implicitly computing a coefficient matrix using low-rank subspace clustering). We then apply a nonnegative feature map (equivalent to implicitly computing an affinity matrix). We then weigh and concatenate the obtained embeddings of each view (equivalent to summing the affinity matrices into a single consensus one). Finally, we perform an SVD followed by a k-Means to obtain a consensus partition (equivalent to performing spectral clustering on the consensus affinity matrix). Novelty comes from extending [fettal2023scalable] to a multi-view setting in an efficient manner through the properties of kernels.
@@ 1 Introduction
The explosive growth of data from diverse sources, such as social media, sensor networks, and online platforms, has led to the emergence of complex high-dimensional datasets. These datasets often contain multiple views of the same underlying data, each capturing different aspects or perspectives. Traditional clustering algorithms, designed for single-view data, face significant challenges in effectively capturing intricate relationships and structures that extend across multiple views. In this context, multi-view clustering, which integrates information from multiple views in order to enhance clustering performance, has emerged as a powerful paradigm. It has been successfully applied to various applications including computer vision [yin2015multi, wang2016iterative], natural language processing [kumar2011co, liu2013multi, boutalbi2022tensor], and bioinformatics [sun2013multi, wang2016multi, sun2015multi, yu2020clustering]. For a more in-depth discussion, the reader can refer to [chao2021survey, yu2025review].
@@
In this work, we are specifically interested in multiview subspace clustering [gao2015multi, sun2021scalable, chen2022efficient, cai2023seeking, chen2025adaptive, zhong2025locality]. By exploring the latent subspaces within each view and leveraging the complementary nature of different views, multi-view subspace clustering algorithms can reveal hidden structures that may not be apparent when considering individual views in isolation. This not only leads to more accurate clustering results, but also enables a deeper understanding of the underlying data. Specifically, multi-view clustering for attributed network data comprising interconnected entities with associated attributes offers a powerful framework for analyzing complex datasets and, for this reason, has gained significant attention in the recent literature [fan2020one2multi, lin2021graph, lin2021multi, pan2021multi, liu2024beyond]. One prominent application of multiview clustering for attributed network data is in social network analysis, since social networks often exhibit rich attributes relating to individuals or groups: demographic information, interests, affiliations, etc. [fan2020one2multi, hu2020open]. Another application is in biological research, where multi-omics data can be integrated using multiple graph convolutional networks [wang2021mogonet]. Finally, recommender systems often rely on multiple sources of information, such as user preferences, item attributes, and social connections [kipf2016semi, wu2019simplifying, shchur2018pitfalls].
@@
The computational complexity of multiview subspace clustering has hindered its widespread adoption in real-world applications, although some efficient approaches have been proposed. These can involve, for example, anchor-based techniques to reduce the size of matrices used in optimization [kang2020large, sun2021scalable]. However, our experiments show that existing methods often struggle with efficiency on large-scale multi-view datasets, and therefore, further computational improvements are needed.
To address the scalability issue of multiview subspace clustering, rather than using anchor-based methods, we decided to take advantage of the properties of kernel feature maps. Indeed, the class of random features is a popular technique for accelerating kernel methods in large-scale problems, recognized by prestigious awards. [rahimi2007random] proposed mapping the the input data to a randomized low-dimensional feature space and applying fast linear methods. [cho2009kernel] introduced positive-definite kernel functions mimicking large neural nets, usable in both shallow and deep architectures. [maji2009max] and [vedaldi2012efficient] developed explicit feature maps for various kernels, significantly reducing train/test times. [pham2013fast] introduced Tensor Sketching, a faster technique to approximate polynomial kernels, and [liu2021random] provided a comprehensive survey on random features, summarizing algorithms, theoretical results and their relationship with deep neural networks. It is worth noting that none of these approaches have been harnessed in the context of clustering, and even less so in the context of multiview clustering, which is the focus of this paper. Our main contributions are:
@@
    Second, we propose an efficient optimization strategy that scales well with large graphs. Using the properties of kernel feature maps, our algorithm significantly reduces computational burden while maintaining state-of-the-art clustering performance. This scalability makes our approach suitable for handling real-world applications, since it is able to scale to datasets with millions of data points.
@@
    Third, we conduct extensive experiments, including statistical significance testing, on real-world benchmark networks of different scales and degrees of overlap, to evaluate our algorithm against state-of-the-art multi-view subspace clustering methods and attributed-network-specific multi-view approaches. Our experiments are entirely reproducible, ensuring transparency and enabling others to validate our results. The entirety of our source code is provided  [^1] https://github.com/chakib401/MvSCK.
The remainder of this paper is structured as follows. In Section 2 we discuss related works. Section 3 introduces the necessary background to develop our approach. In Section 4 we develop our proposed method and explore the algorithm and its complexity. Section 5 contains a detailed description of our experiments. Finally, we give our conclusion in Section 6. A schematic representation of our approach is shown in Figure 1. This paper extends our previous work [fettal2023scalable], where we introduced a scalable attributed-graph subspace clustering method. Here, we generalize the approach to multi-view subspace clustering in a scalable manner by leveraging properties of kernel maps, and provide a more comprehensive experimental evaluation on large-scale network datasets.
@@ 2 Related Work
This section discusses prior work relating to scalable multi-view subspace clustering. We also look at graph multi-view clustering, since we use attributed networks to evaluate performance.
@@ 2.1 Scalable Multi-view Subspace Clustering
In [lin2021multi] Large-Scale Multi-View Subspace Clustering in Linear Time ( LMVSC ) [kang2020large] was introduced. This method, where smaller graphs are created for each view and then integrated to enable spectral clustering on a reduced graph size, derives from the concept of anchor graphs. The Scalable Multi-view Subspace Clustering with Unified Anchors ( SMVSC ) [sun2021scalable] approach integrates anchor learning and graph construction into a unified optimization framework, resulting in a more accurate representation of latent data distribution by learned anchors and a more discriminative clustering structure. In [chen2022efficient], an efficient orthogonal multi-view subspace clustering ( OMSC ) approach was proposed, featuring joint anchor learning, graph construction and partitioning. Recently, [Zhang_Wang_Li_Zhang_Liu_Zhu_Liu_Zhou_Luo_2023] developed a Flexible and Diverse Anchor Graph Fusion for Scalable Multi-view Clustering ( FDAGF ) that introduces a fusion strategy for multi-size anchor graphs.
@@ 2.2 Graph Multi-view Clustering
In [pan2021multi], a multi-view contrastive graph clustering ( MCGC ) method was proposed, designed to learn a consensus graph with contrastive regularization over the learned graph matrix. [lin2021multi] introduced a novel multi-view attributed graph clustering ( MAGC ) framework, which incorporates node attributes and graphs. It is an approach that stands out in three important respects: utilizing graph filtering for smooth node representation, rather than deep neural networks; learning a consensus graph from data to address noise and incompleteness in the original graph; and exploring high-order relations through a flexible regularizer design. The MvAGC model [lin2021graph] is a simple yet effective approach for Multi-view Attributed Graph Clustering. It applies a graph filter to smooth the features without relying on neural network parameter learning. Reduction in computational complexity is achieved through a novel anchor point selection strategy. In addition, a new regularizer is introduced to explore high-order neighborhood information. [fettal2023simultaneous] introduced LMGEC , a generic linear model for performing multi-view attributed graph representation learning and clustering simultaneously. It uses a simple weighing scheme based on inertia and single-hop neighborhood propagation.
3 Motivation
Unlike existing multi-view subspace clustering methods that rely heavily on anchor-based graph reductions or iterative optimization, our approach—MvSCK—achieves scalability through the use of explicit kernel feature maps and kernel summation properties. This formulation enables efficient consensus affinity computation without constructing large intermediate graphs or solving costly iterative subproblems. Furthermore, while prior methods such as LMVSC, SMVSC, or FDAGF approximate scalability by learning reduced anchors, they still incur significant memory and computational overhead as dataset size increases. In contrast, MvSCK factorizes the consensus kernel directly through low-dimensional feature embeddings, yielding linear complexity with respect to data size. This design allows our method to scale seamlessly to millions of nodes while maintaining competitive accuracy.
@@ 4 Preliminaries
In what follows, matrices are denoted in boldface uppercase letters and vectors are denoted in boldface lowercase letters. $[n]$ corresponds to the integer set $\{1,\ldots,n\}$. Tr is the trace operator. $\oplus$ denotes column-wise concatenation. $\|.\|$ refers to the $\ell_{2}$ norm, unless otherwise specified.
@@ 4.1 Subspace Clustering
Given a data matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ and a desired number of clusters $k$, most subspace clustering problems rely on the self-expression property, taking the following form:
$\min_{\mathbf{C}}$ $\|\mathbf{X}-\mathbf{C}\mathbf{X}\|^{2}\,+\,\Omega(\mathbf{C})\quad\text{s.t.}\quad\mathbf{C}\in\mathcal{C}$ (1)
where $\mathbf{C}\in\mathbb{R}^{n\times n}$ is the coefficient matrix, $\Omega$ is a regularization function, and $\mathcal{C}$ is the set of feasible solutions.
In this context, we assume the data lies on a union of linear subspaces. A linear subspace $\mathcal{S}\subseteq\mathbb{R}^{d}$ is a set closed under linear combinations; geometrically, this represents a flat shape (like a line or a plane) passing through the origin. If a data point $\mathbf{x}_{i}$ belongs to subspace $\mathcal{S}$, it can be reconstructed as a linear combination of other points within $\mathcal{S}$.
Once a solution $\mathbf{C}$ is obtained, a subspace affinity graph is constructed, represented by its adjacency matrix $\mathbf{W}$. Typically, this is calculated as:
$\mathbf{W}=\frac{|\mathbf{C}|+|\mathbf{C}^{\top}|}{2}.$ (2)
Finally, a partition of the $k$ groups is obtained by spectral clustering on $\mathbf{W}$. Different values of $\Omega$ and $\mathcal{C}$ correspond to different subspace clustering techniques. For example, it is possible to seek sparse solutions by using the $\ell_{1}$ norm:
$\min_{\mathbf{C}}$ $\|\mathbf{X}-\mathbf{C}\mathbf{X}\|^{2}\,+\,\lambda\|\mathbf{C}\|_{1}\quad\text{s.t.}\quad\text{diag}(\mathbf{C})=0$ (3)
where $\|\cdot\|_{1}$ denotes the element-wise $\ell_{1}$ norm (sum of absolute values) and the constraint $\text{diag}(\mathbf{C})=0$ prevents a point from representing itself.
@@ 4.2 Scalable Subspace Clustering
In [fettal2022subspace, fettal2023scalable, fettal2023boosting], the authors proposed a scalable low-rank subspace clustering model. Specifically, the problem they considered is the following:
$\min_{\mathbf{U}}$ $\|\mathbf{X}-\mathbf{U}\mathbf{U}^{\top}\mathbf{X}\|^{2}\quad\text{s.t.}\quad\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}.$ (4)
To address the problem they use nonnegative kernel maps to create an affinity matrix $\mathbf{W}=\Phi(\mathbf{U})\Phi(\mathbf{U})^{\top}.$ The advantage of computing the affinity matrix in this manner is that spectral partitioning can be obtained by singular values decomposition on $\Phi(\mathbf{U})$ instead of performing eigendecomposition on $\mathbf{W}$, leading to substantial gains in spatial and computational efficiencies.
@@ 4.3 Multi-view Subspace Clustering
$\min_{\mathbf{F},\mathbf{C}_{1},\ldots,\mathbf{C}_{V}}$ $\sum_{v}\|\mathbf{X}_{v}-\mathbf{C}_{v}\mathbf{X}_{v}\|^{2}+\texttt{Tr}(\mathbf{F}^{\top}\mathbf{C}_{v}\mathbf{F})$ (5)
s.t. $\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1},\quad\mathbf{C}_{1},\ldots,\mathbf{C}_{V}\in\mathcal{C},$
where $\mathbf{C}_{v}$ is the coefficient matrix of the $v$-th view and $\mathbf{F}$ is a consensus partition matrix. However, although the block structures in different view-specific coefficient matrices are similar, the magnitudes and signs of the entries in each $C_{v}$ can vary greatly. To solve this problem, [gao2015multi] proposed using affinity matrices $\mathbf{W}_{v}$ rather than coefficient matrices $\mathbf{C}_{v}$. We generalize their formulation by eliminating the specific constraints that the authors used. This results in the following generic multi-view subspace clustering problem:
$\min_{\mathbf{F},\mathbf{C}_{1},..,\mathbf{C}_{V}}$ $\quad\sum_{v}\|\mathbf{X}_{v}-\mathbf{C}_{v}\mathbf{X}_{v}\|^{2}+\texttt{Tr}\left(\mathbf{F}^{\top}\left(\mathbf{I}-\tfrac{|\mathbf{C}_{v}|+|\mathbf{C}_{v}^{\top}|}{2}\right)\mathbf{F}\right)$ (6)
s.t. $\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1},\quad\mathbf{C}_{1},\ldots,\mathbf{C}_{V}\in\mathcal{C}.$
@@
5 Proposed approach
Starting from the multi-view framework presented in problem (6) and, as in (4), we introduce low-rank constraints, also using the same affinity matrix. However, here we consider a symmetrically normalized affinity matrix rather than the unnormalized one. The use of such a matrix can improve the stability, interpretability, performance, and convergence of clustering algorithms (see, for example, [ng2001spectral]). We therefore propose the following objective function:
$\min_{\mathbf{F},\mathbf{U}_{1},\ldots,\mathbf{U}_{V}}$ $\sum_{v}\|\mathbf{X}_{v}-\mathbf{U}_{v}\mathbf{U}_{v}^{\top}\mathbf{X}_{v}\|^{2}+\ \texttt{Tr}(\mathbf{F}^{\top}(\mathbf{I}-\mathbf{W}_{v})\mathbf{F})$ (7)
such that $\mathbf{W}_{v}=\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})\Phi(\mathbf{U}_{v})^{\top}\mathbf{D}_{v}^{-\frac{1}{2}},\qquad\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1}$
where $\mathbf{U}_{v}\in\mathbb{R}^{n\times f}$ is a low-rank matrix of view $v$ ($f<<d$), and $\mathbf{W}_{v}$ is the normalized affinity matrix of view $v$; the degree matrix $D_{v}$ is a diagonal matrix in which each diagonal element represents the degree of the corresponding node $i$ in the graph of view $v$. If we were to address this problem directly, the consensus partition $\mathbf{F}$ would be derived through spectral clustering on